Next Section | Table of Contents


Drilling Strategy and Operations

After arriving on station and lowering the pipe, we planned to triple-APC core the sediment section to refusal, which was estimated to be at ~250 mbsf, to obtain overlapping, and thus complete, coverage for high-resolution paleoenvironmental studies. If time allowed, we then planned to deepen the third hole to 410 mbsf using the XCB. We then hoped to log the open hole using the triple combo and the FMS-sonic tools to provide a quantitative basis for comparison with the multisensor track (MST) data, which could be used to reconstruct a continuous sediment section for the site.

The transit to Site 1202 (proposed Site KS-1) was made in good time with fair seas and favorable currents. The 784-nmi distance was covered in 64.9 hr at an average speed of 12.1 kt. At 0254 hr on the morning of 28 April, the vessel arrived at 24°48.24'N, 122°30.00'E, the coordinates for the drilling location. The crew began lowering thrusters and hydrophones, and at 0515 hr on 28 April, the positioning beacon was deployed.

A standard APC/XCB bottom-hole assembly (BHA), including a lockable float valve (LFV) to allow wireline logging of the deepest hole of the site, was made up. The drill string was tripped to the bottom, and Hole 1202A was spudded at 0810 hr. APC coring continued through Core 195-1202A-9H to a depth of 83.1 mbsf (Table T3) when the APC failed to stroke out. Core 195-1202A-10H fully stroked; however, Cores 11H and 12H did not fully advance. Advance by recovery was used for the two incomplete cores in the hope that the hard layer would be limited in thickness and piston coring would once again become viable. APC refusal was finally accepted when Core 195-1202A-13H at 119.5 mbsf had not only failed to scope, but the core liner failed at the midpoint of the barrel. Core orientation using the Tensor tool was initiated with Core 195-1202A-4H and continued through Core 13H. Temperature measurements were taken on Cores 195-1202A-4H, 7H, 10H, and 13H using the Adara temperature tool. Two of the four temperature measurements were good. The developmental APC-methane tool was deployed on Cores 195-1202A-4H and then on Cores 7H through 13H. All runs were successful in acquiring data. Hole 1202A officially ended with the clearing of the seafloor at 1900 hr on 28 April.

The vessel was offset 15 m to the east, and Hole 1202B was spudded at 1935 hr on 28 April. APC coring continued through Core 195-1202B-13H to a depth of 111.6 mbsf (Table T3) before APC refusal was defined by two successive incomplete strokes on Cores 195-1202B-12H and 13H. Because the ultimate depth objective at this site was 410 mbsf, we decided to cut three XCB cores before terminating the hole to obtain an idea about penetration rates, core recovery, and quality. Coring continued with the XCB through Core 195-1202B-16X to a depth of 140.5 mbsf. The drill string was pulled clear of the seafloor, officially ending Hole 1202B at 0445 hr on 29 April.

The ship was offset 15 m to the east, and Hole 1202C was spudded at 0540 hr on 29 April. APC coring continued through Core 195-1202C-11H to a depth of 97.5 mbsf (Table T1), where APC refusal was encountered as defined by three consecutive incomplete strokes. Core orientation using the Tensor tool was initiated with Core 195-1202C-4H and continued through Core 11H. The drill string was pulled clear of the seafloor, officially ending Hole 1202C at 1215 hr on 29 April.

For the third time, the ship was offset 15 m to the east, and Hole 1202D was spudded with the APC at 1245 hr on 29 April. Recovery of the first core was only 15 cm. APC coring continued in this hole through Core 195-1202D-9H to a depth of 76.2 mbsf (Table T1), when the first core did not achieve full stroke. Coring with the XCB proceeded through Core 195-1202D-32X to a depth of 297.4 mbsf, where a short wiper trip was made to 221.3 mbsf, above an area of poor recovery. The wiper trip was uneventful, and coring continued through Core 195-1202-44X to a depth of 410.0 mbsf, the approved target depth for Site 1202.

In preparation for logging, a wiper trip was initiated at 1930 hr on 30 April and reached the logging depth of 80.0 mbsf without incident. The hole was displaced with 150 barrels of sepiolite logging mud and the bit pulled back to a logging depth of 80.0 mbsf.

The triple combo tool string was made up with the Lamont Doherty Earth Observatory (LDEO) temperature/acceleration/pressure (TAP) tool. A nuclear source was not incorporated because the loss of density data did not outweigh the risk of losing the source in disputed waters under hostile current conditions. Throughout operations at the site, the Kuroshio Current was strong. Heavy pipe vibration was experienced continually, while the currents varied cyclically between 2.6 and nearly 4.0 kt. The first tool string was deployed to a depth of ~215 meters below rig floor. At that point, the logging engineer reported losing all power to the logging tools. After bringing the tool string to the surface, the tools showed several loose connections caused by the current-induced drill sting vibrations. All joints were retightened and taped with duct tape. The tool string was once again deployed inside the drill pipe; however, the winch operator only reached 72 mbsf before losing weight with the logging line, as if setting down on an obstruction. Upon recovery, the connections were once again found to be loose and the lower portion of the TAP tool was missing.

At 0430 hr, the circulating head was rigged up and the coring line was run in the hole to determine if the TAP tool was lodged in the drill string. Results were inconclusive. We decided to abandon further wireline logging efforts because of the intensity of the current-induced drill string vibration. The drill string was pulled clear of the seafloor by 0730 hr. By 1100 hr, preparations were under way to secure and clean the ship for transit into port. The 55-nmi transit to Keelung, Taiwan, was made at reduced speed for an 0815-hr arrival at the pilot station on 2 May. The ship was dockside at 0904 hr, ending Leg 195.

Next Section | Table of Contents